OpenStudy (anonymous):

A cylindrical tank holds 11,550 in3 of water. Iha empties the tank into 25 smaller identical cylinders, each with a radius of 2.71 in. What is the height of the 25 smaller cylinders? Use 3.14 to approximate pi and round the answer to two decimal places.

OpenStudy (anonymous):

i think its 20.03

OpenStudy (jdoe0001):

hmmm how did you get 20.03 anyway?

OpenStudy (jdoe0001):

do you happen to recall the volume of a cylinder's formula?

OpenStudy (welshfella):

the volume of water in the big tank = volume water in 25 of small cylinders so if we let the height of the small ones be h we have: volume of cylinder = 3.14*r^2* h 11,550 = 25 * 3.14*2.71^2 * h h = 11,550 / (25*3.14*2.71^2)

OpenStudy (anonymous):

isnt the formula